OpenStudy (anonymous):

Meh. I don't understand square roots. xD Help, please? I'll type up the problem in the comments.

OpenStudy (anonymous):

Nakim simplified \[3\sqrt{2} + x \sqrt{8x} - 5\sqrt{18x}\] and got \[-10x \sqrt{2x}\] for an answer. Explain what Nakim did wrong, and then simply the expression correctly.

OpenStudy (amistre64):

sqrts are just like factors; you can treat them as an extra variable if need be

OpenStudy (amistre64):

notice that the first one aint got no x in it, so it aint like the others.

OpenStudy (anonymous):

I'm confused. :/

OpenStudy (amistre64):

im sure you are, but i cant read minds so youll have to be more specific in your condition.

OpenStudy (anonymous):

How would I simplify the expressions so I can solve it?

OpenStudy (amistre64):

take out the perfect squares that are inside of them

OpenStudy (amistre64):

can you list the first few perfect squares? up to 5^2 should suffice for this

OpenStudy (anonymous):

Wait, what? D:

OpenStudy (amistre64):

what does "wait what" mean???

OpenStudy (anonymous):

I don't understand what you're asking me to do. :/

OpenStudy (amistre64):

take the first few numbers; 1 - 5 and square them

OpenStudy (anonymous):

whgwertgwhrteg;rgje what first few numbers? x.x

OpenStudy (anonymous):

I seriously don't understand anything you're asking me to do. :/ I don't understand square roots at all.

OpenStudy (amistre64):

apparently you dont understand; first few numbers 1-5 either, or what it means to square them ....

OpenStudy (anonymous):

Apparently not. ._.

OpenStudy (amistre64):

then your not understanding more than just square roots

OpenStudy (anonymous):

Can't you just show me how to simplify one of them and I can use it to help me figure out the other ones? :/

OpenStudy (amistre64):

im trying, but you dont understand

OpenStudy (amistre64):

a square root undoes a square

OpenStudy (amistre64):

what is 3^2 ?

OpenStudy (anonymous):

9?

OpenStudy (amistre64):

yes, and what is 2^2 ?

OpenStudy (anonymous):

4?

OpenStudy (amistre64):

good, these will be all we need for this problem; 4 and 9 are called perfect squares since they can be undone by a square root sqrt(4) = 2 sqrt(9) = 3 does this make sense?

OpenStudy (amistre64):

does 8x have any perfect square factors?

OpenStudy (anonymous):

2 and 4? o.o

OpenStudy (amistre64):

yes :) sqrt(8x) = sqrt(4*2x) = sqrt(4) * sqrt(2x) = 2 sqrt(2x) that is the simplified version of sqrt(8x)

OpenStudy (amistre64):

does 18x have any perfect square factors?

OpenStudy (anonymous):

Oh. The first one is supposed to be 3 sqrt (2x), not 2

OpenStudy (amistre64):

we should keep that in mind after we simplify the others then :)

OpenStudy (anonymous):

and... 18 has 9 and 2. right?

OpenStudy (amistre64):

9 and 2 are correct sqrt(18x) = sqrt(9*2x) = 3sqrt(2x) notice how we can undo a perfect square factor and pull it out

OpenStudy (amistre64):

this leaves us with: \[3\sqrt{2x}-x*2\sqrt{2x}-5*3\sqrt{2x}\]

OpenStudy (anonymous):

So it would be.. 3 sqrt (2x) + 2 sqrt(2x) - 5 * 3sqrt(2x)

OpenStudy (anonymous):

yeah.

OpenStudy (amistre64):

yes, but dont forget the "x" term in the middle

OpenStudy (anonymous):

oops. o.o

OpenStudy (amistre64):

3sqrt(2x) + 2x sqrt(2x) -15sqrt(2x) and the rest is history

OpenStudy (anonymous):

So that means my final answer is.. 2x - 12sqrt(2x)?

OpenStudy (amistre64):

if yo udo it that way, try using this for the notation (2x - 12) sqrt(2x)

OpenStudy (anonymous):

That was my answer before, and my teacher said to "review my final answer." which makes me think that's wrong. :/

OpenStudy (amistre64):

its not wrong; but your teacher might want a different format

OpenStudy (amistre64):

2x sqrt(2x) - 12 sqrt(2x) is another way to write it

OpenStudy (anonymous):

What Nakim did wrong, is he ignored that the second term had a factor of x, and added all the terms together. The real answer would be: 2x - 12sqrt(2x) Response Feedback: Please review your final answer. ^^ that's the response she gave me.

OpenStudy (amistre64):

your notation is bad ... put paranthesis around 2x - 12 to keep it as a single value

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

thanks. n.n

OpenStudy (amistre64):

spose you had: 3n + 2xn -15n 2x -12n is not a proper simplification

OpenStudy (anonymous):

True. I think I see what I did wrong. Thanks for the help. :

OpenStudy (amistre64):

youre welcome